If you pay $20 to enter a game and the winning a prize money of $1000 is 1 out of 200, $500 is 1 out of 100, $200 is 1 out of 50, $100 is 1 out of 20. Do you lose or win in a long run?
Answer by stanbon(75887)

If you pay $20 to enter a game and the winning a prize money of $1000 is 1 out of 200, $500 is 1 out of 100, $200 is 1 out of 50, $100 is 1 out of 20. Do you lose or win in a long run?
----
E(x) = (1/200)*1000 + (1/100)*500 + (1/50)200 + (1/20)*100 - (83/100)20
-----
= (1/200)*1000 + (2/200)*500 + (4/200)200 + (10/200)*100 - (183/200)20
------------------
= (1000+1000+800+1000-3660)/200 = $0.70
--------------
> 0
----
You win.
===========================
Cheers,
Stan H.
============